# Risk-Sensitive Agent Compositions

### v2

@@ Abstract
From software development to robot control, modern agentic systems decompose complex objectives into a sequence of subtasks and choose a set of specialized AI agents to complete them. We formalize agentic workflows as directed acyclic graphs, called agent graphs, where edges represent AI agents and paths correspond to feasible compositions of agents. Real-world deployment requires selecting agent compositions that not only maximize task success but also minimize violations of safety, fairness, and privacy requirements which demands a careful analysis of the low-probability (tail) behaviors of compositions of agents. In this work, we consider risk minimization over the set of feasible agent compositions and seek to minimize the value-at-risk of the loss distribution of the agent composition where the loss quantifies violations of these requirements. We introduce an efficient algorithm which traverses the agent graph and finds a near-optimal composition of agents. It uses a dynamic programming approach to approximate the value-at-risk of agent compositions by exploiting a union bound. Furthermore, we prove that the approximation is near-optimal asymptotically for a broad class of practical loss functions. To evaluate our framework, we consider a suite of video game-like control benchmarks that require composing several agents trained with reinforcement learning and demonstrate our algorithm’s effectiveness in approximating the value-at-risk and identifying the optimal agent composition.
@@ 1 Introduction
Modern agentic workflows orchestrate specialized AI agents to tackle complex tasks demanding a diverse range of skills. Typically, a high-level planning agent decomposes the main task into a sequence of subtasks and selects appropriate worker agents to execute them. These worker agents can be powerful generalist models, such as large language models (LLMs) and vision-language models (VLMs), or control policies trained with reinforcement learning or imitation learning. Such systems have demonstrated considerable success in automating tasks across various domains, including software development [33, 25, 15], complex information retrieval [34], scientific discovery [13], and robot control [17, 9, 32, 35].
In addition to maximizing task success, deploying agentic systems in the real world requires minimizing various forms of risk. The stochastic nature of both environment and agents means that we have to rigorously analyze low-probability but high-consequence tail behaviors that emerge when multiple agents interact. Our work follows a rich body of literature in risk-sensitive planning and control that dates back to the 1970s [14, 31] where the objective is to optimize a risk measure of the loss (or reward) distribution as opposed to the expectation. A risk measure is a function that maps a loss distribution to a real number and takes into account attributes like the variance and the tail of the distribution. Some examples include the value-at-risk (VaR) which corresponds to the tail quantile and conditional value-at-risk (CVaR) which is the expected tail loss. In this work, we consider loss functions that quantify violations of requirements such as safety, fairness, and privacy.
In our formalization, we represent an agentic workflow by an agent graph (example in Figure˜1) which is a directed acyclic graph in which edges correspond to agents (denoted $\pi_{1},\dots,\pi_{4}$) and paths correspond to feasible agent compositions that achieve the overall objective. In Figure˜1, we have two compositions of agents to pick from: $p_{A}\coloneqq\pi_{1}\to\pi_{3}$ and $p_{B}\coloneqq\pi_{2}\to\pi_{4}$. Finally, the risk associated with an agent’s behavior is quantified with a real-valued loss function $L_{i}:T_{i}\to\mathbb{R}$ that takes as input a trace of execution of the corresponding agent $\pi_{i}$.
@@
Example 1: DroneNav. Consider four control agents $\pi_{1},\dots,\pi_{4}$ that have been trained to navigate a drone between the rooms $S\xrightarrow{\pi_{1}}A$, $S\xrightarrow{\pi_{2}}B$, $A\xrightarrow{\pi_{3}}F$, and $B\xrightarrow{\pi_{4}}F$. We define loss functions $L_{i}:T_{i}\to\mathbb{R}$ for each agent that capture safety requirements. They take as input the trajectory from a rollout of a policy and return the negative minimum distance between the obstacles in the rooms and the trajectory. Now, consider the task of navigating a drone from room $S$ to the target room $F$ which can be achieved by two paths: one passing through room $A$ and another through room $B$. This can be represented by the agent graph in Figure˜1 where we have the two compositions of agents to choose from. The loss functions of the compositions of agents $p_{A}$ and $p_{B}$ correspond to the negative minimum distance between the obstacles and the sequence of trajectories of the agents in each composition. In particular, for the composition $p_{A}$, the loss function is denoted $L_{p_{A}}:T_{1}\times T_{3}\to\mathbb{R}$ and is given by $L_{p_{A}}(t_{1},t_{3})\coloneqq\max(L_{1}(t_{1}),L_{3}(t_{3}))$ where $t_{1}$ and $t_{3}$ are trajectories from rollouts of $\pi_{1}$ and $\pi_{3}$ respectively. $L_{p_{B}}$ is symmetrically defined.
@@
Example 2: Information retrieval. Suppose we have access to an LLM agent that we are using in an information retrieval and processing task. In this setting, a natural requirement is to minimize the amount of hallucinated information, i.e., data that is not taken from data sources or the LLM context. We define the loss function to be a metric that quantifies the amount of hallucinated information in the outputs which can be a score produced by an LLM-as-a-Judge. Next, consider the task of constructing an agentic system that takes as input the name of a country and produces a graph of the evolution of literacy rates in the country over a period of five decades. Also, suppose we have two data sources—UNESCO and the World Bank—from which the system needs to fetch relevant pages and then write a Python script to generate the plot. This can be represented by the agent graph in Figure˜1 where $\pi_{1}$ and $\pi_{2}$ represent LLM agents tasked with retrieving relevant text data from the two sources, and $\pi_{3}$ and $\pi_{4}$ represent agents that write the script to produce the plot from the data. Yet again, the loss of the composition of agents is given by $L_{p_{A}}(t_{1},t_{3})\coloneqq\max(L_{1}(t_{1}),L_{3}(t_{3}))$ which now corresponds to the maximum amount of information hallucinated by either of the two agents $\pi_{1}$ or $\pi_{3}$ in their corresponding outputs $t_{1}$ and $t_{3}$.
@@
As noted in these examples, when the loss of each agent quantifies violations of safety, fairness, and privacy requirements, the loss of a composition of agents—the quantity we want to minimize—is the maximum of the losses incurred by each agent. In contrast, existing literature in risk-sensitive planning usually considers the cumulative loss [1, 20] which is the natural choice when the losses represent costs.
@@
Given an agent graph, our objective is to find the path representing an agent composition that minimizes the value-at-risk, denoted $\textnormal{VaR}_{\alpha}$, which is the $(1-\alpha)$-quantile of the loss distribution of the agent composition for a user-defined risk budget $\alpha>0$. In other words, minimizing $\textnormal{VaR}_{\alpha}$ minimizes a constant $\ell^{*}>0$ such that with probability at least $1-\alpha$, each agent in the chosen composition of agents incurs a loss of at most $\ell^{*}$. In the rest of this work, we assume only black-box access to the agents and focus on estimating the value-at-risk through sampling.
@@
A straightforward approach to minimizing the value-at-risk is to consider each feasible agent composition, estimate its tail quantile, and pick the optimal composition. However, this method is inefficient since the graph can have an exponential number of compositions as a function of the number of agents. We overcome this hurdle with an efficient algorithm that scales polynomially in the number of agents and incrementally builds a near-optimal path while traversing the agent graph. At its heart is a quantile estimation method that approximates the value-at-risk of a composition of agents by dividing the risk budget $\alpha$ amongst the agents and applying a union bound. It then finds the optimal allocation of the risk budget over a discretized set of budgets using a dynamic programming algorithm. We prove that this approximation is near-optimal asymptotically when the losses of agents are given by independent random variables. This independence assumption holds when the losses are functions of the agent behavior and do not directly impact task accuracy or performance. Indeed, safety, fairness, and privacy are often properties of the behavior and are orthogonal to task success.
@@
We implement our algorithm and test its performance on a series of reinforcement learning (RL) environments. Here, we look at long-horizon tasks that require composing several agents trained with RL to complete them. We consider two tasks—safety in navigation and fair resource consumption—and design suitable loss functions that capture these requirements. We confirm empirically that our algorithm produces tight approximations of the value-at-risk thereby finding optimal agent compositions.
@@
    We formalize the risk minimization objective for agentic workflows as the problem of finding an agent composition that minimizes the value-at-risk of the loss quantifying violations of safety, fairness, and privacy requirements (Section˜2).
@@
    Then, we introduce an efficient algorithm that approximates this optimization problem with a quantile estimation method using the union bound and dynamic programming. Furthermore, we prove that this approximation is near-optimal asymptotically under independence assumptions. (Section˜3)
@@
    Finally, we evaluate our algorithm on compositional reinforcement learning benchmarks using two tasks: safety-critical navigation and fair resource consumption. We design appropriate loss functions for these tasks and demonstrate that our algorithm successfully identifies optimal agent compositions while providing tight approximations of their value-at-risk (Section˜4).
@@ 1.1 Related work
Our work builds upon a large body of literature on automated planning and control under uncertainty [11]. Traditionally, the planning problem is modeled as a Markov decision process (MDP) over which the expected (un)discounted cumulative loss is minimized using algorithms like value iteration or reinforcement learning. Towards designing risk-aware agents, recent works have instead considered the optimization of risk measures that look at the tail of the loss distribution instead of the expected loss. Ahmadi et al. [1], Kretínský and Meggendorfer [20] study risk minimization for reachability and mean-payoff objectives in MDPs, while Bastani et al. [4], Wang et al. [29], Greenberg et al. [12], Choi et al. [6] develop risk-sensitive reinforcement learning algorithms. These works consider risk measures that fall into the class of coherent risk measures [2] that include the conditional value-at-risk and the entropic value-at-risk. Risk-sensitive objectives have also been considered in the stochastic control literature. Methods for risk-sensitive linear, quadratic, and Gaussian control [31] have been developed along with more recent methods for non-linear model predictive control [24]. We also refer to Wang and Chapman [30] for a comprehensive survey of this topic. However, all of these works stand in contrast with our framework where we want to minimize a risk measure of the maximum incurred loss instead of the cumulative loss.
Related lines of research in robotics include hierarchical reinforcement learning [18, 7, 36, 16, 21] and task and motion planning (TAMP) [19] in which long-horizon control objectives are decomposed into subtasks that are carried out by motion planners or other low-level controllers trained with RL or imitation learning. Similar frameworks have been proposed for automatic generation of agentic workflows to automate software development tasks [25, 33, 15] that sequence LLM-agents to complete smaller subtasks. Agentic systems have also been proposed for the tasks of complex information retrieval [34] and scientific discovery [13]. These methods typically optimize the overall task success probability and do not consider risk minimization.
@@ 2 Risk Minimization Over Agent Graphs
Observe that given a distribution over the inputs $\mathcal{D}(X)$, the agent $f$ induces distributions $\mathcal{D}_{f}(T)$ and $\mathcal{D}_{f}(Y)$ over the set of traces and outputs respectively. We can sample $t\sim\mathcal{D}_{f}(T)$ or $y\sim\mathcal{D}_{f}(Y)$ by first sampling $x\sim\mathcal{D}(X)$, then sampling $(t,y)\sim f(x)$, and dropping one of the two components. Evaluating the loss on the distribution of traces gives us a distribution over the losses the agent incurs. Since we are concerned with risk minimization, we focus on analyzing tail behaviors of the loss distribution. In particular, the quantity we are interested in optimizing is the value-at-risk at level $\alpha\in(0,1)$, denoted $\textnormal{VaR}_{\alpha}$. $\textnormal{VaR}_{\alpha}$ is a risk measure of the loss distribution that corresponds to the $(1-\alpha)$-quantile of the distribution. In other words, a $(1-\alpha)$ fraction of the traces will have a loss smaller than the $\textnormal{VaR}_{\alpha}$. To formally define it, let $Z_{f}\sim\mathcal{D}_{f}(T)$ be a random variable over the traces of the agent $f$. Then we can express
$\textnormal{VaR}_{\alpha}[L(Z_{f})]\coloneqq\textnormal{quantile}(L(Z_{f}),1-\alpha)=\inf\left\{q\in\mathbb{R}:\mathbf{Pr}[L(Z_{f})\leq q]\geq 1-\alpha\right\}$ (1)
@@
Let $G=(V,E,X,T,F,L,s,t,\mathcal{D}_{s})$ be an agent graph and let $\mathcal{P}\subseteq V^{*}$ be the set of directed paths from $s$ to $t$. For a given risk level $\alpha\in(0,1)$ we define the risk minimization objective as the following optimization problem
$\operatorname*{arg\,min}_{p\in\mathcal{P}}\textnormal{VaR}_{\alpha}[L_{p}(Z_{p})].$ (3)
Rephrasing in words, the optimization objective corresponds to finding a path from $s$ to $t$ in the agent graph $G$ that minimizes the value-at-risk of the maximum of the losses incurred by the agents along the path.
3 Finding the path minimizes the value-at-risk
Algorithm BucketedVaR : Efficient algorithm for the RMAG objective

### v3

@@ Abstract
From software development to robot control, modern agentic systems decompose complex objectives into a sequence of subtasks and choose a set of specialized AI agents to complete them. We formalize agentic workflows as directed acyclic graphs, called agent graphs, where edges represent AI agents and paths correspond to feasible compositions of agents. Real-world deployment requires selecting agent compositions that not only maximize task success but also minimize violations of safety, fairness, and privacy requirements which demands a careful analysis of the low-probability (tail) behaviors of compositions of agents. In this work, we consider risk minimization over the set of feasible agent compositions and seek to minimize the value-at-risk and the conditional value-at-risk of the loss distribution of the agent composition where the loss quantifies violations of these requirements. We introduce an efficient algorithm which traverses the agent graph and finds a near-optimal composition of agents. It uses a dynamic programming approach to approximate the value-at-risk of agent compositions by exploiting a union bound. Furthermore, we prove that the approximation is near-optimal asymptotically for a broad class of practical loss functions. We also show how our algorithm can be used to approximate the conditional value-at-risk as a byproduct. To evaluate our framework, we consider a suite of video game-like control benchmarks that require composing several agents trained with reinforcement learning and demonstrate our algorithm’s effectiveness in approximating the value-at-risk and identifying the optimal agent composition.
@@ 1 Introduction
Modern agentic workflows orchestrate specialized AI agents to tackle complex tasks demanding a diverse range of skills. Typically, a high-level planning agent decomposes the main task into a sequence of subtasks and selects appropriate worker agents to execute them. These worker agents can be powerful generalist models, such as large language models (LLMs) and vision-language models (VLMs), or control policies trained with reinforcement learning or imitation learning. Such systems have demonstrated considerable success in automating tasks across various domains, including software development [Zhang et al., 2025a, Niu et al., 2025, Hu et al., 2025], complex information retrieval [Zhang et al., 2025b], scientific discovery [Gridach et al., 2025], and robot control [Ichter et al., 2022, Feng et al., 2025, Yang et al., 2024, Zhao et al., 2025].
In addition to maximizing task success, deploying agentic systems in the real world requires minimizing various forms of risk. The stochastic nature of both environment and agents means that we have to rigorously analyze low-probability but high-consequence tail behaviors that emerge when multiple agents interact. Our work follows a rich body of literature in risk-sensitive planning and control that dates back to the 1970s [Howard and Matheson, 1972, Whittle, 1981] where the objective is to optimize a risk measure of the loss (or reward) distribution as opposed to the expectation. A risk measure is a function that maps a loss distribution to a real number and takes into account attributes like the variance and the tail of the distribution. Some examples include the value-at-risk (VaR) which corresponds to the tail quantile and conditional value-at-risk (CVaR) which is the expected tail loss. In this work, we consider loss functions that quantify violations of requirements such as safety, fairness, and privacy.
In our formalization, we represent an agentic workflow by an agent graph (example in Figure 1) which is a directed acyclic graph in which edges correspond to agents (denoted $\pi_{1},\dots,\pi_{4}$) and paths correspond to feasible agent compositions that achieve the overall objective. In Figure 1, we have two compositions of agents to pick from: $p_{A}\coloneqq\pi_{1}\to\pi_{3}$ and $p_{B}\coloneqq\pi_{2}\to\pi_{4}$. Finally, the risk associated with an agent’s behavior is quantified with a real-valued loss function $L_{i}:T_{i}\to\mathbb{R}$ that takes as input a trace of execution of the corresponding agent $\pi_{i}$.
@@
Example 1: DroneNav. Consider four control agents $\pi_{1},\dots,\pi_{4}$ that have been trained to navigate a drone between the rooms $S\xrightarrow{\pi_{1}}A$, $S\xrightarrow{\pi_{2}}B$, $A\xrightarrow{\pi_{3}}F$, and $B\xrightarrow{\pi_{4}}F$. We define loss functions $L_{i}:T_{i}\to\mathbb{R}$ for each agent that capture safety requirements. They take as input the trajectory from a rollout of a policy and return the negative minimum distance between the obstacles in the rooms and the trajectory. Now, consider the task of navigating a drone from room $S$ to the target room $F$ which can be achieved by two paths: one passing through room $A$ and another through room $B$. This can be represented by the agent graph in Figure 1 where we have the two compositions of agents to choose from. The loss functions of the compositions of agents $p_{A}$ and $p_{B}$ correspond to the negative minimum distance between the obstacles and the sequence of trajectories of the agents in each composition. In particular, for the composition $p_{A}$, the loss function is denoted $L_{p_{A}}:T_{1}\times T_{3}\to\mathbb{R}$ and is given by $L_{p_{A}}(t_{1},t_{3})\coloneqq\max(L_{1}(t_{1}),L_{3}(t_{3}))$ where $t_{1}$ and $t_{3}$ are trajectories from rollouts of $\pi_{1}$ and $\pi_{3}$ respectively. $L_{p_{B}}$ is symmetrically defined.
@@
Example 2: Information retrieval. Suppose we have access to an LLM agent that we are using in an information retrieval and processing task. In this setting, a natural requirement is to minimize the amount of hallucinated information, i.e., data that is not taken from data sources or the LLM context. We define the loss function to be a metric that quantifies the amount of hallucinated information in the outputs which can be a score produced by an LLM-as-a-Judge. Next, consider the task of constructing an agentic system that takes as input the name of a country and produces a graph of the evolution of literacy rates in the country over a period of five decades. Also, suppose we have two data sources—UNESCO and the World Bank—from which the system needs to fetch relevant pages and then write a Python script to generate the plot. This can be represented by the agent graph in Figure 1 where $\pi_{1}$ and $\pi_{2}$ represent LLM agents tasked with retrieving relevant text data from the two sources, and $\pi_{3}$ and $\pi_{4}$ represent agents that write the script to produce the plot from the data. Yet again, the loss of the composition of agents is given by $L_{p_{A}}(t_{1},t_{3})\coloneqq\max(L_{1}(t_{1}),L_{3}(t_{3}))$ which now corresponds to the maximum amount of information hallucinated by either of the two agents $\pi_{1}$ or $\pi_{3}$ in their corresponding outputs $t_{1}$ and $t_{3}$.
@@
As noted in these examples, when the loss of each agent quantifies violations of safety, fairness, and privacy requirements, the loss of a composition of agents—the quantity we want to minimize—is the maximum of the losses incurred by each agent. In contrast, existing literature in risk-sensitive planning usually considers the cumulative loss [Ahmadi et al., 2021, Kretínský and Meggendorfer, 2018] which is the natural choice when the losses represent costs.
@@
Given an agent graph, our objective is to find the path representing an agent composition that minimizes the value-at-risk, denoted $\textnormal{VaR}_{\alpha}$, which is the $(1-\alpha)$-quantile of the loss distribution of the agent composition for a user-defined risk budget $\alpha>0$. In other words, minimizing $\textnormal{VaR}_{\alpha}$ minimizes a constant $\ell^{*}>0$ such that with probability at least $1-\alpha$, each agent in the chosen composition of agents incurs a loss of at most $\ell^{*}$. We also consider another risk measure known as the conditional value-at-risk (also known as expected shortfall), denoted $\textnormal{CVaR}_{\alpha}$, which is defined as the expected loss in the worst $\alpha$ fraction of the loss distribution. In the rest of this work, we assume only black-box access to the agents and focus on estimating these risk measures through sampling.
@@
A straightforward approach to minimizing the value-at-risk is to consider each feasible agent composition, estimate its tail quantile, and pick the optimal composition. However, this method is inefficient since the graph can have an exponential number of compositions as a function of the number of agents. We overcome this hurdle with an efficient algorithm that scales polynomially in the number of agents and incrementally builds a near-optimal path while traversing the agent graph. At its heart is a quantile estimation method that approximates the value-at-risk of a composition of agents by dividing the risk budget $\alpha$ amongst the agents and applying a union bound. It then finds the optimal allocation of the risk budget over a discretized set of budgets using a dynamic programming algorithm. We prove that this approximation is near-optimal asymptotically when the losses of agents are given by independent random variables. This independence assumption holds when the losses are functions of the agent behavior and do not directly impact task accuracy or performance. Indeed, safety, fairness, and privacy are often properties of the behavior and are orthogonal to task success. Further, since the conditional value-at-risk can be approximated as the average of tail quantiles, we show how it can be recovered as a byproduct of our algorithm.
@@
We implement our algorithm and test its performance on a series of reinforcement learning (RL) environments. Here, we look at long-horizon tasks that require composing several agents trained with RL to complete them. We consider two tasks—safety in navigation and fair resource consumption—and design suitable loss functions that capture these requirements. We confirm empirically that our algorithm produces tight approximations of the value-at-risk and the conditional value-at-risk thereby finding optimal agent compositions. Further experiments show that our algorithm is robust to a reasonable amount of correlation between agent losses.
@@
    We formalize the risk minimization objective for agentic workflows as the problem of finding an agent composition that minimizes the value-at-risk of the loss quantifying violations of safety, fairness, and privacy requirements (Section 2).
@@
    Then, we introduce an efficient algorithm that approximates this optimization problem with a quantile estimation method using the union bound and dynamic programming. Furthermore, we prove that this approximation is near-optimal asymptotically under independence assumptions. (Section 3)
@@
    Finally, we evaluate our algorithm on compositional reinforcement learning benchmarks using two tasks: safety-critical navigation and fair resource consumption. We design appropriate loss functions for these tasks and demonstrate that our algorithm successfully identifies optimal agent compositions while providing tight approximations of their risk measures (Section 4).
@@ 1.1 Related work
Our work builds upon a large body of literature on automated planning and control under uncertainty [Ghallab et al., 2016]. Traditionally, the planning problem is modeled as a Markov decision process (MDP) over which the expected (un)discounted cumulative loss is minimized using algorithms like value iteration or reinforcement learning. Towards designing risk-aware agents, recent works have instead considered the optimization of risk measures that look at the tail of the loss distribution instead of the expected loss. Ahmadi et al. [2021], Kretínský and Meggendorfer [2018] study risk minimization for reachability and mean-payoff objectives in MDPs, while Bastani et al. [2022], Wang et al. [2023], Greenberg et al. [2022], Choi et al. [2021] develop risk-sensitive reinforcement learning algorithms. These works consider risk measures that fall into the class of coherent risk measures [Artzner et al., 1999] that include the conditional value-at-risk and the entropic value-at-risk. Risk-sensitive objectives have also been considered in the stochastic control literature. Methods for risk-sensitive linear, quadratic, and Gaussian control [Whittle, 1981] have been developed along with more recent methods for non-linear model predictive control [Nishimura et al., 2020]. We also refer to Wang and Chapman [2022] for a comprehensive survey of this topic. However, all of these works stand in contrast with our framework where we want to minimize a risk measure of the maximum incurred loss instead of the cumulative loss.
Related lines of research in robotics include hierarchical reinforcement learning [Jothimurugan et al., 2021, Dalal et al., 2024, Zhou et al., 2024, Huang et al., 2023, Lin et al., 2024] and task and motion planning (TAMP) [Kaelbling and Lozano-Pérez, 2011] in which long-horizon control objectives are decomposed into subtasks that are carried out by motion planners or other low-level controllers trained with RL or imitation learning. Similar frameworks have been proposed for automatic generation of agentic workflows to automate software development tasks [Niu et al., 2025, Zhang et al., 2025a, Hu et al., 2025] that sequence LLM-agents to complete smaller subtasks. Agentic systems have also been proposed for the tasks of complex information retrieval [Zhang et al., 2025b] and scientific discovery [Gridach et al., 2025]. These methods typically optimize the overall task success probability and do not consider risk minimization.
@@ 2 Risk Minimization Over Agent Graphs
Observe that given a distribution over the inputs $\mathcal{D}(X)$, the agent $f$ induces distributions $\mathcal{D}_{f}(T)$ and $\mathcal{D}_{f}(Y)$ over the set of traces and outputs respectively. We can sample $t\sim\mathcal{D}_{f}(T)$ or $y\sim\mathcal{D}_{f}(Y)$ by first sampling $x\sim\mathcal{D}(X)$, then sampling $(t,y)\sim f(x)$, and dropping one of the two components. Evaluating the loss on the distribution of traces gives us a distribution over the losses the agent incurs. Since we are concerned with risk minimization, we focus on analyzing tail behaviors of the loss distribution. In particular, we are interested in optimizing the following risk measures: the value-at-risk at level $\alpha\in(0,1)$, denoted $\textnormal{VaR}_{\alpha}$ and the conditional value-at-risk, denoted $\textnormal{CVaR}_{\alpha}$. $\textnormal{VaR}_{\alpha}$ corresponds to the $(1-\alpha)$-quantile of the distribution. In other words, a $(1-\alpha)$ fraction of the traces will have a loss smaller than the $\textnormal{VaR}_{\alpha}$. To formally define it, let $Z_{f}\sim\mathcal{D}_{f}(T)$ be a random variable over the traces of the agent $f$. Then we can express
$\textnormal{VaR}_{\alpha}[L(Z_{f})]\coloneqq\textnormal{quantile}(L(Z_{f}),1-\alpha)=\inf\left\{q\in\mathbb{R}:\mathbf{Pr}[L(Z_{f})\leq q]\geq 1-\alpha\right\}.$ (1)
On the other hand, $\textnormal{CVaR}_{\alpha}$, a coherent risk measure Artzner et al. [1999], is more sensitive to the tail loss distribution and is defined as the expected tail loss:
$\textnormal{CVaR}_{\alpha}[L(Z_{f})]\coloneqq\frac{1}{\alpha}\int_{0}^{\alpha}\textnormal{VaR}_{\gamma}[L(Z_{f})]d\gamma=\mathbb{E}[L(Z_{f})\mid L(Z_{f})\geq\textnormal{VaR}_{\alpha}[L(Z_{f})]].$ (2)
@@
Let $G=(V,E,X,T,F,L,s,t,\mathcal{D}_{s})$ be an agent graph and let $\mathcal{P}\subseteq V^{*}$ be the set of directed paths from $s$ to $t$. For a given risk level $\alpha\in(0,1)$ and risk measure $\rho\in\{\textnormal{VaR}_{\alpha},\textnormal{CVaR}_{\alpha}\}$ we define the risk minimization objective as the following optimization problem
$\operatorname*{arg\,min}_{p\in\mathcal{P}}\rho[L_{p}(Z_{p})].$ (4)
Rephrasing in words, the optimization objective corresponds to finding a path from $s$ to $t$ in the agent graph $G$ that minimizes the risk measure of the maximum of the losses incurred by the agents along the path.
3 Finding the path minimizing risk
Algorithm BucketedVaR : Efficient algorithm for the RMAG-VaR objective